can you show me how to do inqualities using addition and subtraction. for example: -6(4x+3)=-8(10+x)


Answer by checkley77(12844)

-6(4x+3)=-8(10+x)
-24x-18=-80-8x
-24x+8x=-80+18
-16x=-62
x=-62/-16
x=3.875
Proof:
-6(4*3.875+3)=-8(10+3.875)
-6(15.5+3)=-8*13.875
-6*18.5=-111
-111=-111